Selected T-account balances for Fields Company are shown below as of January 31, 2019; adjusting entries have already been posted. The firm uses a calendar-year accounting period but prepares monthly adjustments.

Supplies (A)
Jan.31 Bal 800
Supplies Expense (E)
Jan.31 Bal 960
Prepaid Insurance (A)
Jan.31 Bal 574
Insurance Expense (E)
Jan.31 Bal 82
Wages Payable (L)
500 Jan.31 Bal
Wages Expense (E)
Jan.31 Bal 3,200
Truck (A)
Jan.31 Bal 8,700
Accumulated Depreciation-Truck (XA)
2,610 Jan.31 Bal

(a) If the amount in Supplies Expense represents the January 31 adjustment for the supplies used in January, and $620 worth of supplies were purchased during January, what was the January 1 beginning balance of Supplies?

(b) The amount in the Insurance Expense account represents the adjustment made at January 31 for January insurance expense. If the original insurance premium was for one year, what was the amount of the premium?

What date did the insurance policy start?


(c) If we assume that no beginning balance existed in Wages Payable or Wages Expense on January 1, how much cash was paid as wages during January?


(d) If the truck has a useful life of five years, what is the monthly amount of depreciation expense?


How many months has Fields owned the truck?
_______ months
Answer #1

(a)
Supplies Beginning Balance = $800+960-620 = $1140

(b)
Amount of Premium = $82 x 12 = $984

Balance Months left = $584/82 = 7 months
i.e. 5 months insurance have expired
So Date of Policy start = Sep 1

(c)
Cash Paid for wages expense = $3200-500 = $2700

(d)
Monthly Depreciation expense = $8700 / 60 = $145

No of months truck owned = $2610 / 145 = 18 months

If you have any query, kindly comment with your query and please mark thumbs up.